# Patch-wise Structural Loss for Time Series Forecasting

### v1

@@ Abstract
Time-series forecasting has gained significant attention in machine learning due to its crucial role in various domains. However, most existing forecasting models rely heavily on point-wise loss functions like Mean Square Error, which treat each time step independently and neglect the structural dependencies inherent in time series data, making it challenging to capture complex temporal patterns accurately. To address these challenges, we propose a novel Patch-wise Structural (PS) loss, designed to enhance structural alignment by comparing time series at the patch level. Through leveraging local statistical properties, such as correlation, variance, and mean, PS loss captures nuanced structural discrepancies overlooked by traditional point-wise losses. Furthermore, it integrates seamlessly with point-wise loss, simultaneously addressing local structural inconsistencies and individual time-step errors. PS loss establishes a novel benchmark for accurately modeling complex time series data and provides a new perspective on time series loss function design. Extensive experiments demonstrate that PS loss significantly improves the performance of state-of-the-art models across diverse real-world datasets. The data and code are publicly available at: https://github.com/Dilfiraa/PS_Loss.
@@ 1 Introduction
Time series forecasting plays a crucial role across various real-world domains, including traffic (Kong et al., 2024; Long et al., 2024), weather (Lam et al., 2023; Wu et al., 2023b), and finance (Huang et al., 2024a), where accurate forecasting is essential for informed decision-making. Recent advancements in deep learning have developed models that capture complex dependencies and intricate patterns in time series data, significantly improving forecasting accuracy (Zhou et al., 2021, 2022; Sun et al., 2024a; Yu et al., 2024; Sun et al., 2024b). These models aim to produce multi-step predictions based on historical data that closely follow the trajectory of future series.
@@ 2.1 Time Series Forecasting Models
In recent years, numerous deep forecasting models have been proposed to capture complex dependencies in time series data by leveraging the powerful representation capabilities of neural networks (Kim et al., 2024). Based on their architecture, these models can be categorized into CNN, Transformer, MLP, and LLM-based approaches. CNN-based models excel at modeling local temporal relationships (Wu et al., 2023a; Luo & Wang, 2024; Cheng et al., 2024; Li et al., 2024; Liu et al., 2022). In contrast, Transformer-based models have gained significant attention for their ability to model long-range dependencies through self-attention mechanisms (Vaswani et al., 2017; Nie et al., 2023; Liu et al., 2024a; Chen et al., 2024b; Lin et al., 2024b; Zhou et al., 2022, 2021). Additionally, MLP-based models offer efficient alternatives that balance simplicity with performance (Zeng et al., 2023; Wang et al., 2024b; Yi et al., 2024; Lin et al., 2024a; Ekambaram et al., 2023). Recently, LLM-based models have emerged as a promising direction, leveraging the capabilities of large language models to effectively capture complex temporal dependencies (Zhou et al., 2023; Liu et al., 2024b). These models use point-wise MSE as their loss function, treating each time step independently, which limits their ability to capture the structural nuances of time series data and may hinder forecasting performance.
@@ 2.2 Loss Functions for Time Series Forecasting
Recent efforts to address the limitations of MSE have given rise to alternative loss functions (Le Guen & Thome, 2019; Cuturi & Blondel, 2017; Xiong et al., 2024; Qi et al., 2024), which can be broadly categorized into shape-focused losses and dependency-focused losses. Shape-focused losses aim to capture structural similarities between the ground truth and predictions by explicitly addressing shape misalignment. For instance, Dynamic Time Warping (DTW)-based methods such as Soft-DTW (Cuturi & Blondel, 2017) and DILATE (Le Guen & Thome, 2019) align sequences under temporal distortions. While effective at improving shape alignment, their high computational complexity limits scalability. TILDE-Q (Lee et al., 2022), on the other hand, introduces transformation invariance, making it robust to amplitude shifts, phase shifts, and scaling differences, thus focusing on shape similarity. Dependency-focused losses focus on capturing temporal dependencies within the forecast window. FreDF (Wang et al., 2024a) bypasses label correlation complexity by learning to forecast in the frequency domain. Despite these advancements, most prior methods rely on global comparisons of entire time series, overlooking crucial localized structural details. In contrast, our proposed approach incorporates patch-wise statistical properties into the loss function, enabling a more granular structural measurement of the data and providing a fundamentally new perspective on time series loss design.
@@ 3.2 Overview
We first generate predictions $\hat{\mathbf{Y}}$ using a backbone model, forming the basis for subsequent patch-level comparisons. Following this, FAP segments both the ground truth and prediction into patches based on the dominant frequency of the ground truth, ensuring that the patches capture recurring patterns and coherent structural information. Once the series is segmented, the PS loss evaluates the similarity between patches by integrating correlation, variance, and mean losses, enabling a comprehensive assessment of local structural alignment across the time series. To further refine the optimization process, GDW dynamically adjusts the weights of the three losses based on their gradients, ensuring balanced contributions from each component and leading to more effective optimization and consistent model training.
@@ 3.3 Fourier-based Adaptive Patching
In order to adaptively determine the patch length, we introduce the Fourier-based Adaptive Patching (FPA) technique. By analyzing the amplitude spectrum of the ground truth via the Fast Fourier Transform (FFT), we identify its dominant periodicity:
$\mathbf{A}=\operatorname{Amp}\left(\mathcal{FFT}(\mathbf{Y})\right),\quad f=\underset{f_{*}\in\{1,\ldots,\lceil T/2\rceil\}}{\arg\max}(\mathbf{A}),$ (2)
where, $\operatorname{Amp}(\cdot)$ calculates the amplitude values and $f$ denotes the dominant frequency. The period $p=\lceil T/f\rceil$ serves as an initial patch length candidate, ensuring the patches reflect recurring structural patterns. However, dominant periodicity alone may produce overly large patches, missing finer structure nuances. To balance granularity and coherence, the final patch length is determined as:
$P=\min\left(\left\lceil\frac{p}{2}\right\rceil,\delta\right),\quad p=\left\lceil\frac{T}{f}\right\rceil,$ (3)
@@
With patch length $P$ and stride $S=\lceil P/2\rceil$, the series is segmented into patches $\hat{\mathbf{Y}}_{p}\in\mathbb{R}^{P\times N}$ and $\mathbf{Y}_{p}\in\mathbb{R}^{P\times N}$, where $N=\lfloor(T-P)/S\rfloor+1$ is the total number of patches. The $i$-th patch of $\mathbf{Y}$ is defined as $\mathbf{Y}_{p}^{(i)}=\{y_{0}^{(i)},y_{1}^{(i)},\ldots,y_{P-1}^{(i)}\}$, where $y_{j}^{(i)}$ refers to the $y_{i\times S+j}$ in the unpatched time series.
@@ 3.4 Patch-wise Structural Loss
Building on the adaptive patching framework, we propose the Patch-wise Structural (PS) loss to enhance time series forecasting by capturing and aligning localized statistical properties within each patch. PS loss focuses on three key metrics: correlation, variance, and mean - to offer a deeper structural comparison.
@@
where $\mathbf{\mu}^{(i)}$ and $\mathbf{\sigma}^{(i)}$ represent the mean and standard deviation of the $i$-th patch, respectively, and $\rho(\cdot,\cdot)$ denotes the PCC. Minimizing $\mathcal{L}_{Corr}$ encourages the model to align its predictions with the ground truth’s directionality and pattern trends. This alignment helps the model capture underlying pattern relationships, which are critical for time series forecasting tasks that require more than just point-wise accuracy.
@@
where $\alpha$, $\beta$, and $\gamma$ balance the contributions of the individual components. This integration provides a comprehensive structural alignment by leveraging the strengths of each component loss.
@@ 4.1 Experimental Setup
Models iTransformer (2024a) PatchTST (2023) TimeMixer (2024b) DLinear (2023) TimseNet (2023a)
@@
Datasets. We conduct our experiments on seven real-world multivariate time series datasets, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Weather, ECL, and Exchange. For data preprocessing, we follow the standard protocol used by backbone studies. The detailed experimental setup is described in Appendix A.
Backbones. We select five state-of-the-art (SOTA) time series forecasting models with diverse architectures to comprehensively evaluate PS loss function. Specifically, we include Transformer-based models: iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023); MLP-based models: DLinear (Zeng et al., 2023), TimeMixer (Wang et al., 2024b); CNN-based model: TimesNet (Wu et al., 2023a).
Implementation Details. We use the official implementations of each backbone from their respective repositories for our experiments. For fair evaluation, when integrating PS loss to enhance the backbone model’s performance, we follow their original experimental and hyperparameter settings, while only tuning the PS loss weight $\lambda$ and the patch length threshold $\delta$. All the experiments are implemented by PyTorch on a single NVIDIA RTX 3090 24GB GPU.
@@ 4.2 Main Results
Table 2: LLM-based long-term multivariate forecasting results. The table reports MSE and MAE for different forecasting lengths $T\in\{96,192,336,720\}$. The input sequence length is consistent with the backbone setting, $L=336$ for OFA and $L=672$ for AutoTimes. The better results are highlighted in bold.
@@ 4.4 Comparison with Other Loss Functions
We evaluate the proposed PS loss against advanced loss functions. TILDE-Q (Lee et al., 2022) emphasizes shape similarity using transformation-invariant loss terms, while FreDF (Wang et al., 2024a) bypasses label correlation by comparing time series in the frequency domain. The results are shown in Table 3. PS loss achieves the lowest MSE and MAE in most cases across various datasets and forecasting horizons. This is due to its ability to effectively measure localized structural similarity by evaluating correlation, variance, and mean at the patch level, which enables a more precise alignment between the ground truth and predictions.
@@ 4.5 Ablation Study
To evaluate the contribution of each component within the proposed PS loss, we conducted ablation experiments on the ETTh1 and weather datasets using DLinear and iTransformer as the backbone models. The results are presented in Table 4, with additional results provided in Appendix H.
@@
Dynamic weighting. To evaluate the importance of dynamic weighting, we replaced the adaptive weighting scheme with a fixed weight of 1 for all three loss terms. This static weighting strategy results in performance degradation, demonstrating that dynamic weighting effectively balances each loss term and improves optimization and forecasting.
Table 5: Zero-shot forecasting results on ETT datasets. The table reports MSE and MAE values for backbone models trained with and without PS loss for zero-shot forecasting. The forecasting length is 192.
@@
Loss Function MSE loss + PS loss MSE loss + PS loss MSE loss + PS loss
@@ 4.7 Forecasting Visualization
We visualize samples from the Weather dataset using iTransformer as the backbone model. As shown in Figure 3, PS loss significantly improves alignment with the ground truth compared to MSE loss: (a) its predictions closely follow the trajectory of the ground truth, preserving overall trends and patterns. (b) it effectively captures amplitude fluctuations, reflecting the localized variability. (c) it reduces biases at the overall level, ensuring better alignment with the central tendency of the ground truth. These results highlight the strengths of PS loss in capturing localized structural patterns. More visualizations are provided in Appendix G.
@@ Appendix B Backbone Models
We select seven state-of-the-art (SOTA) models from different architectures as backbones. These include Transformer-based models: iTransformer (Liu et al., 2024a) and PatchTST (Nie et al., 2023); MLP-based models: DLinear (Zeng et al., 2023) and TimeMixer (Wang et al., 2024b); CNN-based model: TimesNet (Wu et al., 2023a); LLM-based models: OFA (Zhou et al., 2023) and AutoTimes (Liu et al., 2024b). We use the official implementations of each backbone from their respective repositories and follow their optimal hyperparameter settings for our experiments.
@@ Appendix C Related Work
Patching Mechanism in Time Series Forecasting. In recent years, patch-based approaches have gained significant traction in time series forecasting, offering substantial improvements in prediction performance (Zhang & Yan, 2023; Chen et al., 2024b; Tang & Zhang, 2024). By segmenting the input time series into patches, these methods enhance locality and capture comprehensive semantic information that point-level techniques often overlook (Nie et al., 2023).

### v2

@@ Abstract
Time-series forecasting has gained significant attention in machine learning due to its crucial role in various domains. However, most existing forecasting models rely heavily on point-wise loss functions like Mean Squared Error, which treat each time step independently and neglect the structural dependencies inherent in time series data, making it challenging to capture complex temporal patterns accurately. To address these challenges, we propose a novel Patch-wise Structural (PS) loss, designed to enhance structural alignment by comparing time series at the patch level. Through leveraging local statistical properties, such as correlation, variance, and mean, PS loss captures nuanced structural discrepancies overlooked by traditional point-wise losses. Furthermore, it integrates seamlessly with point-wise loss, simultaneously addressing local structural inconsistencies and individual time-step errors. PS loss establishes a novel benchmark for accurately modeling complex time series data and provides a new perspective on time series loss function design. Extensive experiments demonstrate that PS loss significantly improves the performance of state-of-the-art models across diverse real-world datasets. The data and code are publicly available at: https://github.com/Dilfiraa/PS_Loss.
@@ 1 Introduction
Time series forecasting plays a crucial role across various real-world domains, including traffic (Kong et al., 2024; Long et al., 2024), weather (Lam et al., 2023; Wu et al., 2023b), and finance (Huang et al., 2024a), where accurate forecasting is essential for informed decision-making. Recent advancements in deep learning have developed models that capture complex dependencies and intricate patterns in time series data, significantly improving forecasting accuracy (Zhou et al., 2021, 2022; Sun et al., 2024a; Yu et al., 2024b; Sun et al., 2024b; Yu et al., 2024a; Qiu et al., 2025b; Sun et al., 2025). These models aim to produce multi-step predictions based on historical data that closely follow the trajectory of future series.
@@ 2.1 Time Series Forecasting Models
In recent years, numerous deep forecasting models have been proposed to capture complex dependencies in time series data by leveraging the powerful representation capabilities of neural networks (Kim et al., 2025; Shao et al., 2024; Qiu et al., 2024, 2025a). Based on their architecture, these models can be categorized into CNN, Transformer, MLP, and LLM-based approaches. CNN-based models excel at modeling local temporal relationships (Wu et al., 2023a; Luo & Wang, 2024; Liu et al., 2022). In contrast, Transformer-based models have gained significant attention for their ability to model long-range dependencies through self-attention mechanisms (Nie et al., 2023; Liu et al., 2024a; Chen et al., 2024b; Lin et al., 2024b; Zhou et al., 2022; Yu et al., 2023). Additionally, MLP-based models offer efficient alternatives that balance simplicity with performance (Zeng et al., 2023; Wang et al., 2024; Lin et al., 2024a; Ekambaram et al., 2023; Jia et al., 2025). Recently, LLM-based models have emerged as a promising direction, leveraging the capabilities of large language models to effectively capture complex temporal dependencies (Zhou et al., 2023; Liu et al., 2024b; Jin et al., 2023; Niu et al., 2025). These models use point-wise MSE as their loss function, treating each time step independently, which limits their ability to capture the structural nuances of time series data and may hinder forecasting performance.
@@ 2.2 Loss Functions for Time Series Forecasting
Recent efforts to address the limitations of MSE have given rise to alternative loss functions (Le Guen & Thome, 2019; Cuturi & Blondel, 2017; Lee et al., 2022; Wang et al., 2025), which can be broadly categorized into shape-focused losses and dependency-focused losses. Shape-focused losses aim to capture structural similarities between the ground truth and predictions by explicitly addressing shape misalignment. For instance, Dynamic Time Warping (DTW)-based methods such as Soft-DTW (Cuturi & Blondel, 2017) and DILATE (Le Guen & Thome, 2019) align sequences under temporal distortions. While effective at improving shape alignment, their high computational complexity limits scalability. TILDE-Q (Lee et al., 2022), on the other hand, introduces transformation invariance, making it robust to amplitude shifts, phase shifts, and scaling differences, thus focusing on shape similarity. Dependency-focused losses focus on capturing temporal dependencies within the forecast window. FreDF (Wang et al., 2025) bypasses label correlation complexity by learning to forecast in the frequency domain. Despite these advancements, most prior methods rely on global comparisons of entire time series, overlooking crucial localized structural details. In contrast, our proposed approach incorporates patch-wise statistical properties into the loss function, enabling a more granular structural measurement of the data and providing a fundamentally new perspective on time series loss design.
@@ 3.2 Overview
We first generate the prediction $\hat{\mathbf{Y}}$ using a backbone model, forming the basis for subsequent patch-level comparisons. Following this, FAP segments both the ground truth and prediction into patches based on the dominant frequency of the ground truth, ensuring that the patches capture recurring patterns and coherent structural information. Once the series is segmented, the PS loss evaluates the similarity between patches by integrating correlation, variance, and mean losses, enabling a comprehensive assessment of local structural alignment across the time series. To further refine the optimization process, GDW dynamically adjusts the weights of the three losses based on their gradients, ensuring balanced contributions from each component and leading to more effective optimization and consistent model training.
@@ 3.3 Fourier-based Adaptive Patching
In order to adaptively determine the patch length, we introduce the Fourier-based Adaptive Patching (FAP) technique. By analyzing the amplitude spectrum of the ground truth via the Fast Fourier Transform (FFT), we identify its dominant periodicity:
$\mathbf{A}=\operatorname{Amp}\left(\mathcal{FFT}(\mathbf{Y})\right),\quad f=\underset{f_{*}\in\{1,\ldots,\lfloor T/2\rfloor\}}{\arg\max}(\mathbf{A}),$ (2)
where, $\operatorname{Amp}(\cdot)$ calculates the amplitude values and $f$ denotes the dominant frequency. The period $p=\lfloor T/f\rfloor$ serves as an initial patch length candidate, ensuring the patches reflect recurring structural patterns. However, dominant periodicity alone may produce overly large patches, missing finer structure nuances. To balance granularity and coherence, the final patch length is determined as:
$P=\min\left(\left\lfloor\frac{p}{2}\right\rfloor,\delta\right),\quad p=\left\lfloor\frac{T}{f}\right\rfloor,$ (3)
@@
With patch length $P$ and stride $S=\lfloor P/2\rfloor$, the series is segmented into patches $\hat{\mathbf{Y}}_{p}\in\mathbb{R}^{P\times N}$ and $\mathbf{Y}_{p}\in\mathbb{R}^{P\times N}$, where $N=\lfloor(T-P)/S\rfloor+1$ is the total number of patches. The $i$-th patch of $\mathbf{Y}$ is defined as ${Y}_{p}^{(i)}=\{y_{0}^{(i)},y_{1}^{(i)},\ldots,y_{P-1}^{(i)}\}$, where $y_{j}^{(i)}$ refers to the $y_{i\times S+j}$ in the unpatched time series.
@@ 3.4 Patch-wise Structural Loss
Building on the adaptive patching framework, we propose the Patch-wise Structural (PS) loss to enhance time series forecasting by capturing and aligning localized statistical properties within each patch. PS loss focuses on three key metrics, namely correlation, variance, and mean, to offer a deeper structural comparison.
@@
where $\mathbf{\mu}^{(i)}$ and $\mathbf{\sigma}^{(i)}$ represent the mean and standard deviation of the $i$-th patch, respectively, and $\rho(\cdot,\cdot)$ denotes the PCC. Minimizing $\mathcal{L}_{Corr}$ encourages the model to align its predictions with the ground truth’s directionality and trends. This alignment helps the model capture underlying pattern relationships, which are critical for time series forecasting tasks that require more than just point-wise accuracy.
@@
where $\alpha$, $\beta$, and $\gamma$ balance the contributions of the individual loss components. This integration provides a comprehensive structural alignment by leveraging the complementary strengths of each loss.
@@ 4.1 Experimental Setup
Models iTransformer (2024a) PatchTST (2023) TimeMixer (2024) DLinear (2023) TimseNet (2023a)
@@
Datasets. We conduct our experiments on seven real-world multivariate time series datasets, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Weather, ECL, and Exchange. For data preprocessing, we follow the standard protocol used by backbone studies. The detailed description of the datasets is provided in Appendix A.
Backbones. We select five state-of-the-art (SOTA) time series forecasting models with diverse architectures to comprehensively evaluate PS loss function. Specifically, we include Transformer-based models: iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023); MLP-based models: DLinear (Zeng et al., 2023), TimeMixer (Wang et al., 2024); CNN-based model: TimesNet (Wu et al., 2023a).
Implementation Details. We use the official implementations of each backbone from their respective repositories for our experiments. For fair evaluation, when integrating PS loss to enhance the backbone model’s performance, we follow their original experimental and hyperparameter settings, while only tuning the PS loss weight $\lambda$ and the patch length threshold $\delta$. The experiments are implemented using PyTorch and executed on NVIDIA RTX 3090 24GB GPU.
@@ 4.2 Main Results
Table 2: LLM-based long-term multivariate forecasting results. The table reports MSE and MAE for different forecasting lengths $T\in\{96,192,336,720\}$. The input sequence length is consistent with the backbone setting. The better results are highlighted in bold.
@@ 4.4 Comparison with Other Loss Functions
We evaluate the proposed PS loss against advanced loss functions. TILDE-Q (Lee et al., 2022) emphasizes shape similarity using transformation-invariant loss terms, while FreDF (Wang et al., 2025) bypasses label correlation by comparing time series in the frequency domain. The results are shown in Table 3. PS loss achieves the lowest MSE and MAE in most cases across various datasets and forecasting horizons. This is due to its ability to effectively measure localized structural similarity by evaluating correlation, variance, and mean at the patch level, which enables a more precise alignment between the ground truth and predictions.
@@ 4.5 Ablation Study
To evaluate the contribution of each component within the proposed PS loss, we conducted ablation experiments on the ETTh1 and weather datasets using DLinear as the backbone models. The results are presented in Table 4, with additional results provided in Appendix I.
@@
Dynamic weighting. To evaluate the importance of dynamic weighting, we replaced the adaptive weighting with a fixed weight of 1.0 for all three loss terms. This static weighting strategy results in performance degradation, demonstrating that dynamic weighting effectively balances each loss term and improves optimization and forecasting.
Table 5: Zero-shot forecasting results on ETT datasets. The table reports MSE and MAE values for backbone models trained with and without PS for zero-shot forecasting. The forecasting length is 192.
@@
Loss Function MSE + PS MSE + PS MSE + PS
@@ 4.7 Forecasting Visualization
We visualize samples from the Weather dataset using iTransformer as the backbone model. As shown in Figure 3, PS loss significantly improves alignment with the ground truth compared to MSE loss: (a) its predictions closely follow the trajectory of the ground truth, preserving overall trends and patterns. (b) it effectively captures amplitude fluctuations, reflecting the localized variability. (c) it reduces biases at the overall level, ensuring better alignment with the central tendency of the ground truth. These results highlight the strengths of PS loss in capturing localized structural patterns. More visualizations are provided in Appendix H.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62376194, 61925602, U23B2049 and China Scholarship Council Grant 202406250137.
@@ Appendix B Backbone Models
We select seven state-of-the-art (SOTA) models from different architectures as backbones. These include Transformer-based models: iTransformer (Liu et al., 2024a) and PatchTST (Nie et al., 2023); MLP-based models: DLinear (Zeng et al., 2023) and TimeMixer (Wang et al., 2024); CNN-based model: TimesNet (Wu et al., 2023a); LLM-based models: OFA (Zhou et al., 2023) and AutoTimes (Liu et al., 2024b). We use the official implementations of each backbone from their respective repositories and follow their optimal hyperparameter settings for our experiments.
@@ Appendix C Related Work
Patching Mechanism in Time Series Forecasting. In recent years, patch-based approaches have gained significant traction in time series forecasting, offering substantial improvements in prediction performance (Zhang & Yan, 2023; Chen et al., 2024b). By segmenting the input time series into patches, these methods enhance locality and capture comprehensive semantic information that point-level techniques often overlook (Nie et al., 2023; Zhang & Yan, 2023).